OpenStudy (anonymous):

Calculus Help!! Find the Taylor polynomials centred at 0 for n = 0,1,2 for the function f(x) = arcsin(x). Find a bound on the error for each of these on the interval x\[\in\] [-0.5,0.5].

OpenStudy (anonymous):

\[P _{0}=0\] \[P _{1}=x\] P _{2}=x

OpenStudy (anonymous):

but i am not sure how to do the bound error!!

OpenStudy (anonymous):

\[x \in [-0.5,0.5].\]

OpenStudy (anonymous):

use reminder term in Lagrange form calculatded at max of your interval

OpenStudy (anonymous):

i knew but i am not sure what should i put for the M \[\left| R _{n}(x) \right|=M \left| x-a \right|^{n+1}/(n+1)!\]

OpenStudy (anonymous):

the value of the 4th derivative at the point 0.5

OpenStudy (anonymous):

sry third, in your case

OpenStudy (anonymous):

n = 0,1,2

OpenStudy (anonymous):

i still don't really get it!!

OpenStudy (anonymous):

ok, wait

OpenStudy (anonymous):

\[P(x) = f(0) + f \prime(0)x + f \prime \prime(0)x ^{2}/2! + f \prime \prime \prime(0,5)0.5^{3}/3!\] the last term is you max error

OpenStudy (anonymous):

just find the derivatives till order 3

OpenStudy (anonymous):

ok now?

OpenStudy (anonymous):

yea,thank you!!